rxcklesskaisher:

Lisa has to cut her grandma's grass this weekend and wants to know exactly how much area she will be cutting. Calculate the area of the polygon. Be sure to show all your work and explain your answer.

rxcklesskaisher:

1 attachment
eatahgood:

@TETSXPREME

eatahgood:

i am not so good with explanations

rxcklesskaisher:

@Isry ??

rxcklesskaisher:

im not good at typings of math

eatahgood:

if anything i am that type to just give you the answer

eatahgood:

but i can get in trouble

rxcklesskaisher:

i have to explain it by typing

TETSXPREME:

Do you know the formula?

QuestionCoveBot:

Sorry, this is something I struggle a little in. I will try to help. Maybe @darkknight can help you.

rxcklesskaisher:

L*W*H

TETSXPREME:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @rxcklesskaisher L*W*H \(\color{#0cbb34}{\text{End of Quote}}\) yess

darkknight:

break it up into shapes you can work with, such as 2 triangles and a rectangle, then add the areas up

rxcklesskaisher:

thtats what i struggle with idk what to add its too much numbers.

darkknight:

 break it up into something like this

QuestionCoveBot:

 This would be the polygon broken into different shapes.

rxcklesskaisher:

i know the area of the first one though

darkknight:

area for triangle = 1/2 B * H area for rectangle = L * W just p lu gin numbers andsolve, add up the areas

rxcklesskaisher:

i only got hep with the first one.

AZ:

We finished this question two days ago https://questioncove.com/study#/updates/605f56216247741977d5700

rxcklesskaisher:

yes thats why im looking for the question

AZ:

Did you find the question?

rxcklesskaisher:

no

AZ:

What is your question? What are you looking for?

rxcklesskaisher:

the question we did

AZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ Write down everything we did To calculate the area of the polygon, we have to split it into two trapezoids. The area of the first trapezoid is ... show all your work here The area of the second trapezoid is ... show all your work here The total area of the polygon is area of first trapezoid + area of second trapezoid which would be _____ \(\color{#0cbb34}{\text{End of Quote}}\)

rxcklesskaisher:

ok

AZ:

Does that help?

rxcklesskaisher:

yes thats what i needed

rxcklesskaisher:

wait i need the areas

rxcklesskaisher:

i guess not

AZ:

Go through the question and you can find all the work we did to get the area as well as the area https://questioncove.com/study#/updates/605f56216247741977d5700

rxcklesskaisher:

again??!?!?